OpenStudy (anonymous):

Find the slope of the line passing through (-3,-4) (3,-2)

OpenStudy (anonymous):

Slope is given by: \[\large Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\[(x_1, y_1) = (-3, -4)\] \[(x_2, y_2) = (3, -2)\] Plug in these values and tell me what did you get as Slope...

OpenStudy (anonymous):

Ok So I plotted these on my graphing paper and I get a slope of 3?

OpenStudy (anonymous):

How did you get Slope = 3 ??

OpenStudy (anonymous):

\[\large Slope = \frac{-2 -(-4)}{3 - (-3)} \implies \frac{-2 + 4}{3 + 3} = ??\]

OpenStudy (anonymous):

I am also having trouble with 4x + 8y = 7 3x + 4y = 6 I am getting (5/3, 2/8)

OpenStudy (anonymous):

Do you want to find the slope for these ??

OpenStudy (anonymous):

No I just had to solve

OpenStudy (anonymous):

Just multiply first equation by 3 and second equation by 4 and tell me what do you get..

OpenStudy (anonymous):

I just re did it 4x+8y=7 3x+4y=6 -6x-8y=-12 4x+8y=7 -2x=-5 x=5/2 15/2+4y=12/2 4y=-3/2 y=-3/8